Determine the slope of the line shown at the right.
Graph = (0, 5) & (-4, 0)
find the slope of the line.
m = ?

Answer by Alan3354(69443)

Determine the slope of the line shown at the right.
Graph = (0, 5) & (-4, 0)
------------
m = (y2-y1)/(x2-x1)
m = (5-0)/(0-(-4))
m = 5/4